OpenStudy (anonymous):

If one of the roots of the equation x^2+kx-12=0 is 4, what is the value of k? a. -1 c. 1 b. 0 d. 3

OpenStudy (anonymous):

c

OpenStudy (anonymous):

(x-3)(x+4) = x^2 + x - 12

jhonyy9 (jhonyy9):

16+4k-12=0 4k+4=0 4k=-4 k=-4/4 =-1 a. is true